if |a-b| = |a| = |b| then, what is the angle between a and b?

Grade:11

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

There must be more conditions given, if not

a,b can be zero vectors and angles possible are infinite

But that is trivial, kindly attach the question's complete image

Regards
Akanksha
3 years ago
We are given |a-b| = |a| = |b|
 
a-b = sqrt(a^2 + b^2 – 2abcostheta) = a
 
a^2 = a^2 + a^2 – 2a^2costheta
 
2a^2costheta = a^2
 
costheta = ½
 
Hence theta = 60 deg